# Summon a Demon and Bind it: A Grounded Theory of LLM Red Teaming in the Wild

### v1

@@ 1 Introduction
In late 2022, the confluence of improved large language model (LLM) output and broad easy access to these models in a chat format led to an explosion of interest in LLMs. This included the discovery that, though creative use of the inputs given to them, models could be induced to give output that seemed to violate norms or rules – either implicit ones, or explicit ones state by the models or model creators themselves. We present a study of this activity, undertaken at this unique moment where the technology collides with broader society. The core data of the study is 28 in-depth interviews with people performing this adversairal activity.
@@
The term “Red teaming" is well-defined in its arena of origin, military exercise. Various attempts exist to define the term in relation to information security, also [45, 25]; most agree that is it a semi-structued adversarial exercise with clear objectives: to expose vulnerabilities in products or systems. However, red teaming as it relates to LLMs is a new practice. It is quixotic to lift a precise definition from one theatre of operations and expect it to fit an emerging human activity in a different context mediated using different tools and have differently defined objectives. In this paper, we take an evidence-based approach to examining how people in fact adapt and conduct red teaming and related activities in against a novel target: large language models. Our research questions were:
@@ 2 Background
Human interactions with machine learning models are no exception. Attacking these models in order to understand their weaknesses has become so well-organised and productive that many large technology corporations have dedicate teams who try to make models fail in specific ways; their work is integrated into production processes and their advice is solicited by national governments [27].
These targeted technologies have been often accompanied by a reasonably high barrier to entry. Much specialised knowledge is required to e.g. circumvent manufacturer controls on a mobile phone. In cases where little knowledge is required, the impacts have been relatively low (e.g. a foreign address can be validated by an overconstrained system).
@@
Some people refer to the action of adversarially probing a machine learning model as “red teaming". This term has military heritage, from war games, where those roleplaying the opponents are the “red team". This phrase has also been co-opted in information security in general, and then into machine learning and now also generation AI probing. The literature includes multiple definitiones. There detailed guides to how to perform the exercise in the information technology context [45]. While there are multiple definitions of red teaming, consensus exists that red teaming requires a target goal, rather than being open-ended [25]. Finally, game-theoretic framings of the activity are also given [32]. Rather than constrain our study to any of these strict definitions, we instead considered “LLM red teaming" as a subset of the adversarial activity that people conduct with these models. This choice enables a broader view of the overall activity that subsumes LLM red teaming, especially when conducted outside of these focused corporate “red team" units; thus “red teaming in the wild".
@@
There are also efforts to enumerate and characterise LLM vulnerabilities, mapping out the territory of LLM security. The OWASP Top 10 for LLM [48] [^2] https://llmtop10.com/ highlights the top ten vulnerabilities in application security with regard to LLM. This involves some discussion and definition of what a vulnerability even is in the LLM context, and then through community has selected ten risks to be aware of, including example exploit scenarios. The AI Risk and Vulnerability Alliance (ARVA) [^3] https://avidml.org/arva/ maintains a database of AI vulnerabilities and incidents; these are not limited to LLM holes, but cover all of AI. The goal here is to catalogue vulnerabilities. ACL SIGSEC [^4] https://sig.llmsecurity.net/ is the special interest group of the association for computational linguistics that focuses on language processing and language modelling security, collecting and sharing research in the area. Finally, the US National Institute of Standards and Technology has a generative AI work group whose efforts include understanding risks presented by LLMs. [^5] https://www.nist.gov/video/nist-generative-ai-public-working-group-virtual-tour These efforts each serve to catalog and sometimes even characterise the results of building attacks on LLMs, but do not provide understanding into how and why such attacks are constructed.
@@
Work on optimizing some kinds of attacks is advancing. Liu et al. [31] conducted an empirical study of jailbreaking. They collected 78 prompts from the website jailbreakchat.com and classified them into 10 categories of jailbreak patterns. The authors found that the prompts could consistently evade the restrictions in 40 use-case scenarios for ChatGPT. The taxanomy presented by Liu et al. is similar to the taxonomy we present in the current paper (and the categories are comparable), but based on a different type of data; namely the analysis of concrete prompts from a single source of the internet, rather than the underlying motivations or strategies of the prompt creators. Approaches to building automatic jailbreaking tools have transcended universal jailbreaks which work across many different targets [28], intensified to the point where two emerged simultaneously with the same name [50, 29], and the focus is now on attempt efficiency [6].
@@ 3.2 Participant population
We recruited 28 participants. We were not able to mention per-participant job titles or organisations without deanonymising interviewees, so we instead list employers and job titles directly here, out of order. Job titles held by the population include: analyst, artist, assistant professor, associate professor, computer programmer, distinguished engineer, game designer, head of developer experience, inventory at a weed farm, machine learning engineer, not working, penetration tester, phd student, policy research program manager, research manager, research scientist, senior principal research manager, senior research scientist, senior scientist, software developer, software engineer, software generalist, staff prompt engineer, startup founder, and student. The population’s host organizations included Microsoft, Google, Robust Intelligence, Scale AI, a weed farm, Center for Human-Compatible AI, University College Dublin, University of Toronto, and the Hebrew University of Jerusalem.
Note the male gender skew in participants. This is in part due to the options yield by our sampling strategy, and in part due to interviewee preference. While we approached twelve non-male practictioners, only four agreed to an interview.
@@ 3.3 Interview structure
We used an interview approach between semi-structured and open-ended, where we consider the difference to be that the participants were not explicitly made aware of the aims of our study at the beginning (as they would be in a semi-structured approach), but we exlained this during the debriefing, which often led to more discussion topics. We also allowed a high degree of flexibility in order and wording of many of the questions, which is closer to an open-ended interview approach [4]. The interviews were conducted via video call by one or two of the authors, and all interviews were recorded so the authors could watch them again.
@@ 3.4 Data analysis
The tags were clustered into 16 categories and one uncagetorized group (a full overview is provided in the appendix in Table 3). Through the axial coding we started to identify the typical categories of a grounded theory: causal conditions (what factors cause the core phenomenon), strategies (actions taken in response to the core phenomenon), contextual and intervening conditions (broad and specific situational factors that influence the strategies) and consequences (outcomes from using the strategies) [10].
@@
This quote is not from our data, but from the title of an article on qualitative data and generalization [9]. The goal of a grounded theory is to discover a comprehensive theory grounded thoroughly in the views of the participants – not to make statistical predictions or to verify hypotheses. Indeed, the work of building a grounded theory generally goes before hypothesis construction. Therefore, we explicitly abstain from indicating numbers of participants during the report of the findings. First, it is not possible to meaningfully quantify exactly how many participants ‘thought’ or ‘felt’ something – and it does not make sense to measure it. If one participant felt strongly concerned about the impact of language models on the future of society, and another participant was slightly concerned, does that mean that one and a half participants were concerned? This kind of information does not exist in an inherently metric space. Second, when the purpose is discovery and mapping, outliers are as interesting as the mean: “Human activities contain their own means of generalisation that cannot be reduced to extraneous criteria (numbers of observations, duration of fieldwork, sample size, etc.)” [9]. Even if only a single practitioner of an activity does it a certain way, that is still a way that activity can be performed. Quantitative approaches are therefore of reduced impact here and one must be careful around choosing when and when not to apply them.
@@ 4.1 The Core activity: Red Teaming in the Wild
We annotated 61 different highlights in the data with the tag “Naming of the activity”. By order of magnitude, the follwing were the terms used when we asked participants what they would call the activity (in parentheses, number of participants who used this term): Prompt engineering (6), red teaming (4), Prompt injection (4), gaming (3), [prompt] hacking (3), jailbreaking (2), and exploitation of models (2). Participants frequently used more than one term to name the activity. In addition to that, at least 18 other unique terms were used, from the more colorful such as scrying, alchemy, and spellcasting (three different participants), to the more formal, like adversarial prompting/input/testing and fairness evaluation (also three different participants). One participant declined to describe the activity with words and instead had created an image (Figure 2).
@@ 4.1.1 Characteristic #1: The activity is limit-seeking in nature
During the interviews, several of the participants’ described uses of the model for more intended purposes, i.e., not trying to “break” the models or circumvent their filters. [^8] We note here, that none of the participants claimed to be exclusively adversarial users, and that the sections of the interviews when they described more casual uses have generally been left out of our analyses (or tagged as separate categories). However, there is often not a clear delimitation between intended and adversarial use, and all interactions with the model lead to increased experiential knowledge, relevant to either intent. Such uses included (but were not exclusive to) personal educational purposes (like establishing a vocabulary around some problem that they could then investigate further), rubber ducking, playing games, finding the perfect recipe for crunchy sweet potato fries, and (somewhat tongue in cheek) as an oracle for predicting reality. Such uses are not the focus of this study. For these uses, a succcessful interaction would save the user time (P03), or the generated output would be useful or effective (P10), and these criteria were not present in the evaluation criteria for red teaming activities (Section 4.4.3).
@@
In the interest of defining the core phenomenon, we tagged participants’ uses of metaphors for their adversarial interactions with langugage models. This helps us understand how participants make sense of the model and their own role. In Table 4.1.4 we report the prevailing metaphors.
The most frequently used metaphor is that of a fortress. The second most frequent metaphor is that of the model as an object in space, that can be pushed around and back it into a corner. These metaphors and, to some degree, the metaphor of the model as material share the characteristic of exploring boundaries and limits and potentially crossing them. The fortress-metaphor also enforces, pun inteded, the connection to red teaming, where the goal is to adopt an adversarial approach to find the (security) holes in some system [35].
@@ 4.1.4 Characteristic #4: It is a team effort
Model as captive/servant “I’ll force it to correct whatever it has done” (P08)
@@ 4.2 Causal Conditions: Motivations and goals
In this section, we detail the different motivations and goals our participants had for performing red teaming activities. We use the term goal to describe a future-directed event, something that can be worked towards acheiving and is measurable. We use the term motivation to describe a past-directed accumulation of experiences that drives one to want to achieve a specific goal.
@@ 4.2.1 Motivations
“often it’s more exploratory. A lot of times, especially for probing or red teaming activities, I’m just curious, or I come up with a question like, ‘would it do this?’. And then it’s a sort of a trance state of trial and error of trying to get it to do the right thing or trying different variations. […] I’ll get trapped for an hour or two, doing something I didn’t even really mean to be trying out” (P07)
@@
The exploration of a novel technology is inherently fun to a lot of people, and this tinkering play was generally open-ended (although a more goal-orinted, gaming aspect was also mentioned by some participants, e.g., “now, it has become a game to try to jailbreak, as they say, to break it, and to make it generate anything you want” (P15)). We hypothesize that the tinkering nature of the activity might explain why we found significantly fewer women than men in the purposive sampling, since women have repeatedly been shown to express less interest in tinkering with technology than men [3]. We leave this as a subject for future research.
@@
Some of our participants were employed in a professional red team at a company, and some were researchers in security, safety and natural language processing. Obviously, these participants had a professional motivation or ‘monetary incentives’ (P21). A professional motivation may also include a desire to ‘stay on top of things’ and improve one’s own hirability: “my external motivation, if I want to have a chance of keep working in natural language processing, I better start understanding what is going on here quick” (P02), “it’s a banal answer. I want somebody to hire me to do this. (P07).
We inquired explicityly about other extrinsic motivatins, to which many participants mentioned the social credit for adding valuable content to a discourse, or for getting social media engagement: “I had a viral tweet where I asked ChatGPT to rewrite Baby Got Back in the style of the Canterbury Tales, right? And 100,000 people saw that or something”. [^10] Participant number not disclosed for anonymity.
@@ 4.3 Context
“basically, the whole thing is maybe at most like a dozen people on Twitter that are just active and mess around with this kind of stuff. And you just end up in various group chats, or just participate in this, someone might post something weird and interesting, and you’ll kind of, either adopt their ideas or bounce ideas off each other. And it just ends up being exactly like improv, where you mess around with the same idea over and over” (P19).
The community provides inspiration, knowledge, and ‘community heuristics’ (P01), as well as encouragement in terms of engagement. The community is not gathered on one specific website, but is distrbuted over several different platforms; Twitter, Reddit, and various Discord and Slack channels. Even TikTok was mentioned as a source of sharing prompt injection attacks. There exists in fact an invitation-only Slack populated solely by professional red teamers, independent of employer. Community is an essential influence on red teaming strategies. Only one participant (P10) described once getting one piece of negative feedback for their online content: “by publicizing what the jailbreaks were, I was making it like a new passion. And that was bad. […] Like, I wish you hadn’t done this, wish you hadn’t written this”.
@@
Knowledge management is an important ‘intervening condition’ in the sense that succesful red teaming should ideally involve some degree of rigor or structure [35]. Some participants had a structured approach of managing different evaluations:
@@
There exist no public standards for knowledge or library management, and most of the community’s knowledge is shared in screenshots and scattered around the internet – which is why some of our participants had made a point of gathering examples and annotating them in a blog post or other repository. The lack of structured and shared matrics, standards, and libraries makes the process of acquiring comprehensive knowledge of this activity cumbersome at best.

### v2

@@ 1 Introduction
In late 2022, the confluence of improved large language model (LLM) output and broad easy access to these models in a chat format led to an explosion of interest in LLMs. This included the discovery that, though creative use of the inputs given to them, models could be induced to give output that seemed to violate norms or rules – either implicit ones, or explicit ones state by the models or model creators themselves. We present a study of this activity, undertaken at this unique moment where the technology collides with broader society. The core data of the study is 28 in-depth interviews with people performing this adversarial activity.
@@
The term “Red teaming" is well-defined in its arena of origin, military exercise. Various attempts exist to define the term in relation to information security, also [45, 25]; most agree that is it a semi-structured adversarial exercise with clear objectives: to expose vulnerabilities in products or systems. However, red teaming as it relates to LLMs is a new practice. It is quixotic to lift a precise definition from one theater of operations and expect it to fit an emerging human activity in a different context mediated using different tools and have differently defined objectives. In this paper, we take an evidence-based approach to examining how people in fact adapt and conduct red teaming and related activities in against a novel target: large language models. Our research questions were:
@@ 2 Background
Human interactions with machine learning models are no exception. Attacking these models in order to understand their weaknesses has become so well-organized and productive that many large technology corporations have dedicate teams who try to make models fail in specific ways; their work is integrated into production processes and their advice is solicited by national governments [27].
These targeted technologies have been often accompanied by a reasonably high barrier to entry. Much specialized knowledge is required to e.g. circumvent manufacturer controls on a mobile phone. In cases where little knowledge is required, the impacts have been relatively low (e.g. a foreign address can be validated by an overconstrained system).
@@
Some people refer to the action of adversarially probing a machine learning model as “red teaming". This term has military heritage, from war games, where those roleplaying the opponents are the “red team". This phrase has also been co-opted in information security in general, and then into machine learning and now also generation AI probing. The literature includes multiple definitions. There detailed guides to how to perform the exercise in the information technology context [45]. While there are multiple definitions of red teaming, consensus exists that red teaming requires a target goal, rather than being open-ended [25]. Finally, game-theoretic framings of the activity are also given [32]. Rather than constrain our study to any of these strict definitions, we instead considered “LLM red teaming" as a subset of the adversarial activity that people conduct with these models. This choice enables a broader view of the overall activity that subsumes LLM red teaming, especially when conducted outside of these focused corporate “red team" units; thus “red teaming in the wild".
@@
There are also efforts to enumerate and characterize LLM vulnerabilities, mapping out the territory of LLM security. The OWASP Top 10 for LLM [48] [^2] https://llmtop10.com/ highlights the top ten vulnerabilities in application security with regard to LLM. This involves some discussion and definition of what a vulnerability even is in the LLM context, and then through community has selected ten risks to be aware of, including example exploit scenarios. The AI Risk and Vulnerability Alliance (ARVA) [^3] https://avidml.org/arva/ maintains a database of AI vulnerabilities and incidents; these are not limited to LLM holes, but cover all of AI. The goal here is to catalog vulnerabilities. ACL SIGSEC [^4] https://sig.llmsecurity.net/ is the special interest group of the association for computational linguistics that focuses on language processing and language modeling security, collecting and sharing research in the area. Finally, the US National Institute of Standards and Technology has a generative AI work group whose efforts include understanding risks presented by LLMs. [^5] https://www.nist.gov/video/nist-generative-ai-public-working-group-virtual-tour These efforts each serve to catalog and sometimes even characterize the results of building attacks on LLMs, but do not provide understanding into how and why such attacks are constructed.
@@
Work on optimizing some kinds of attacks is advancing. Liu et al. [31] conducted an empirical study of jailbreaking. They collected 78 prompts from the website jailbreakchat.com and classified them into 10 categories of jailbreak patterns. The authors found that the prompts could consistently evade the restrictions in 40 use-case scenarios for ChatGPT. The taxonomy presented by Liu et al. is similar to the taxonomy we present in the current paper (and the categories are comparable), but based on a different type of data; namely the analysis of concrete prompts from a single source of the internet, rather than the underlying motivations or strategies of the prompt creators. Approaches to building automatic jailbreaking tools have transcended universal jailbreaks which work across many different targets [28], intensified to the point where two emerged simultaneously with the same name [50, 29], and the focus is now on attempt efficiency [6].
@@ 3.2 Participant population
We recruited 28 participants. We were not able to mention per-participant job titles or organizations without de-anonymizing interviewees, so we instead list employers and job titles directly here, out of order. Job titles held by the population include: analyst, artist, assistant professor, associate professor, computer programmer, distinguished engineer, game designer, head of developer experience, inventory at a weed farm, machine learning engineer, not working, penetration tester, phd student, policy research program manager, research manager, research scientist, senior principal research manager, senior research scientist, senior scientist, software developer, software engineer, software generalist, staff prompt engineer, startup founder, and student. The population’s host organizations included Microsoft, Google, Robust Intelligence, Scale AI, a weed farm, Center for Human-Compatible AI, University College Dublin, University of Toronto, and the Hebrew University of Jerusalem.
Note the male gender skew in participants. This is in part due to the options yield by our sampling strategy, and in part due to interviewee preference. While we approached twelve non-male practitioners, only four agreed to an interview.
@@ 3.3 Interview structure
We used an interview approach between semi-structured and open-ended, where we consider the difference to be that the participants were not explicitly made aware of the aims of our study at the beginning (as they would be in a semi-structured approach), but we explained this during the debriefing, which often led to more discussion topics. We also allowed a high degree of flexibility in order and wording of many of the questions, which is closer to an open-ended interview approach [4]. The interviews were conducted via video call by one or two of the authors, and all interviews were recorded so the authors could watch them again.
@@ 3.4 Data analysis
The tags were clustered into 16 categories and one uncategorized group (a full overview is provided in the appendix in Table 3). Through the axial coding we started to identify the typical categories of a grounded theory: causal conditions (what factors cause the core phenomenon), strategies (actions taken in response to the core phenomenon), contextual and intervening conditions (broad and specific situational factors that influence the strategies) and consequences (outcomes from using the strategies) [10].
@@
This quote is not from our data, but from the title of an article on qualitative data and generalization [9]. The goal of a grounded theory is to discover a comprehensive theory grounded thoroughly in the views of the participants – not to make statistical predictions or to verify hypotheses. Indeed, the work of building a grounded theory generally goes before hypothesis construction. Therefore, we explicitly abstain from indicating numbers of participants during the report of the findings. First, it is not possible to meaningfully quantify exactly how many participants ‘thought’ or ‘felt’ something – and it does not make sense to measure it. If one participant felt strongly concerned about the impact of language models on the future of society, and another participant was slightly concerned, does that mean that one and a half participants were concerned? This kind of information does not exist in an inherently metric space. Second, when the purpose is discovery and mapping, outliers are as interesting as the mean: “Human activities contain their own means of generalization that cannot be reduced to extraneous criteria (numbers of observations, duration of fieldwork, sample size, etc.)” [9]. Even if only a single practitioner of an activity does it a certain way, that is still a way that activity can be performed. Quantitative approaches are therefore of reduced impact here and one must be careful around choosing when and when not to apply them.
@@ 4.1 The Core activity: Red Teaming in the Wild
We annotated 61 different highlights in the data with the tag “Naming of the activity”. By order of magnitude, the following were the terms used when we asked participants what they would call the activity (in parentheses, number of participants who used this term): Prompt engineering (6), red teaming (4), Prompt injection (4), gaming (3), [prompt] hacking (3), jailbreaking (2), and exploitation of models (2). Participants frequently used more than one term to name the activity. In addition to that, at least 18 other unique terms were used, from the more colorful such as scrying, alchemy, and spellcasting (three different participants), to the more formal, like adversarial prompting/input/testing and fairness evaluation (also three different participants). One participant declined to describe the activity with words and instead had created an image (Figure 2).
@@ 4.1.1 Characteristic #1: The activity is limit-seeking in nature
During the interviews, several of the participants’ described uses of the model for more intended purposes, i.e., not trying to “break” the models or circumvent their filters. [^8] We note here, that none of the participants claimed to be exclusively adversarial users, and that the sections of the interviews when they described more casual uses have generally been left out of our analyses (or tagged as separate categories). However, there is often not a clear delimitation between intended and adversarial use, and all interactions with the model lead to increased experiential knowledge, relevant to either intent. Such uses included (but were not exclusive to) personal educational purposes (like establishing a vocabulary around some problem that they could then investigate further), rubber ducking, playing games, finding the perfect recipe for crunchy sweet potato fries, and (somewhat tongue in cheek) as an oracle for predicting reality. Such uses are not the focus of this study. For these uses, a successful interaction would save the user time (P03), or the generated output would be useful or effective (P10), and these criteria were not present in the evaluation criteria for red teaming activities (Section 4.4.3).
@@
In the interest of defining the core phenomenon, we tagged participants’ uses of metaphors for their adversarial interactions with language models. This helps us understand how participants make sense of the model and their own role. In Table 4.1.4 we report the prevailing metaphors.
The most frequently used metaphor is that of a fortress. The second most frequent metaphor is that of the model as an object in space, that can be pushed around and back it into a corner. These metaphors and, to some degree, the metaphor of the model as material share the characteristic of exploring boundaries and limits and potentially crossing them. The fortress-metaphor also enforces, pun intended, the connection to red teaming, where the goal is to adopt an adversarial approach to find the (security) holes in some system [35].
@@ 4.1.4 Characteristic #4: It is a team effort
Model as captive/servant “I’$l$ force it to correct whatever it has done” (P08)
@@ 4.2 Causal Conditions: Motivations and goals
In this section, we detail the different motivations and goals our participants had for performing red teaming activities. We use the term goal to describe a future-directed event, something that can be worked towards achieving and is measurable. We use the term motivation to describe a past-directed accumulation of experiences that drives one to want to achieve a specific goal.
@@ 4.2.1 Motivations
“often it’s more exploratory. A lot of times, especially for probing or red teaming activities, I’m just curious, or I come up with a question like, ‘would it do this?’. And then it’s a sort of a trance state of trial and error of trying to get it to do the right thing or trying different variations. […] I’$l$ get trapped for an hour or two, doing something I didn’t even really mean to be trying out” (P07)
@@
The exploration of a novel technology is inherently fun to a lot of people, and this tinkering play was generally open-ended (although a more goal-oriented, gaming aspect was also mentioned by some participants, e.g., “now, it has become a game to try to jailbreak, as they say, to break it, and to make it generate anything you want” (P15)). We hypothesize that the tinkering nature of the activity might explain why we found significantly fewer women than men in the purposive sampling, since women have repeatedly been shown to express less interest in tinkering with technology than men [3]. We leave this as a subject for future research.
@@
Some of our participants were employed in a professional red team at a company, and some were researchers in security, safety and natural language processing. Obviously, these participants had a professional motivation or ‘monetary incentives’ (P21). A professional motivation may also include a desire to ‘stay on top of things’ and improve one’s own hireability: “my external motivation, if I want to have a chance of keep working in natural language processing, I better start understanding what is going on here quick” (P02), “it’s a banal answer. I want somebody to hire me to do this. (P07).
We inquired explicitly about other extrinsic motivations, to which many participants mentioned the social credit for adding valuable content to a discourse, or for getting social media engagement: “I had a viral tweet where I asked ChatGPT to rewrite Baby Got Back in the style of the Canterbury Tales, right? And 100,000 people saw that or something”. [^10] Participant number not disclosed for anonymity.
@@ 4.3 Context
“basically, the whole thing is maybe at most like a dozen people on Twitter that are just active and mess around with this kind of stuff. And you just end up in various group chats, or just participate in this, someone might post something weird and interesting, and you’$l$ kind of, either adopt their ideas or bounce ideas off each other. And it just ends up being exactly like improv, where you mess around with the same idea over and over” (P19).
The community provides inspiration, knowledge, and ‘community heuristics’ (P01), as well as encouragement in terms of engagement. The community is not gathered on one specific website, but is distributed over several different platforms; Twitter, Reddit, and various Discord and Slack channels. There exists in fact an invitation-only Slack populated solely by professional red teamers, independent of employer. Even TikTok was mentioned as a source of sharing prompt injection attacks. Community is an essential influence on red teaming strategies. Only one participant (P10) described once getting one piece of negative feedback for their online content: “by publicizing what the jailbreaks were, I was making it like a new passion. And that was bad. […] Like, I wish you hadn’t done this, wish you hadn’t written this”.
@@
Knowledge management is an important ‘intervening condition’ in the sense that successful red teaming should ideally involve some degree of rigor or structure [35]. Some participants had a structured approach of managing different evaluations:
@@
There exist no public standards for knowledge or library management, and most of the community’s knowledge is shared in screenshots and scattered around the internet – which is why some of our participants had made a point of gathering examples and annotating them in a blog post or other repository. The lack of structured and shared metrics, standards, and libraries makes the process of acquiring comprehensive knowledge of this activity cumbersome at best.